Problem 4. Scaling (a) Use Kleibers Law to determine the basal metabolic rate of a typical adult Norwegian rat, of mass 800 grams, given that typical 65 kg human adults have a basal metabolic rate of about 2,000 (food) Calories/day. Joule is 0.2388 physics calories [call, and 1 food Calorie is 1 kilocalorie [kcal]). (b) Given that the Norwegian rat is a true hibernator, with metabolic rate dropping to 1/3 of its basal rate during hibernation, how much weight will it lose during 5 months of hibernation? (Assume that the rats fat yields 3.8 x 104 J/gl of energy).
Answer #1

a) Kleiber's law: Basal Metabolic Rate = k x (mass).75

Thus 2000 food cal/day = k x (65).75

solving for k = 2000/22.89 = 87.367 cal/kg.75

Thus for Norwegien rat C = 87.367 x (.8).75 = 73.9 food cal/day

b) hibernating MR = 73.9/3 = 24.635 food cal/day

5 months = 5 x 30 = 150 days

Thus, required food cal = 150 x 24.635 =7390.5 food cal = 3695.25 kcal = 3695.25 x 1000 x 4.186 J

this energy is provided by fat only, so the weight of fat consumed = (3695.25 x 1000 x 4.186)/(3.8 x 104) = 407 grams